OpenStudy (clobber02):

2 - (-2) to the power of -2

OpenStudy (clobber02):

@ganeshie8

OpenStudy (anonymous):

is this what you mean:\[2-(-2)^2\]

OpenStudy (clobber02):

no

OpenStudy (clobber02):

\[2-(-2)^{-2}\]

OpenStudy (anonymous):

when you have something on the power of negative, it will become a fraction:\[2-(-\frac{ 1 }{ 4 })^2\]

OpenStudy (clobber02):

k now what?

OpenStudy (anonymous):

square it

OpenStudy (clobber02):

?

OpenStudy (clobber02):

square what?

OpenStudy (anonymous):

actually \[2-(\frac{ 1 }{ 4 })^1\] because you already used the negative with -1.

OpenStudy (anonymous):

\[2-\frac{ 1 }{ 4 }\]

OpenStudy (clobber02):

1.75 or\[\frac{ 3 }{ 4 }\]

OpenStudy (clobber02):

i dont know

OpenStudy (anonymous):

\[\frac{ 8 }{ 4 }-\frac{ 1 }{ 4 }\]

OpenStudy (clobber02):

\[\frac{ 7 }{ 4 }\]

OpenStudy (anonymous):

yep!

OpenStudy (clobber02):

thank u can u help with some more

OpenStudy (anonymous):

sure : )

OpenStudy (anonymous):

anything on the power is equal to 1 \[2^0=1\]

OpenStudy (clobber02):

ok

OpenStudy (anonymous):

so now have: \[(1-1+5^2-0)^{-1}\times5\]

OpenStudy (clobber02):

(25)

OpenStudy (clobber02):

1/4 x 5?

OpenStudy (anonymous):

yep so you have \[(25)^{-1}=?\]

OpenStudy (clobber02):

1/4

OpenStudy (anonymous):

um no....\[\frac{ 1 }{ 25 }\]

OpenStudy (clobber02):

how

OpenStudy (anonymous):

when something is one the power of -1, the numerator and denominator switch if you have 25/1, the power of -1 would be 1/25

OpenStudy (clobber02):

ooooohhhhhh i get it now so now its 1/25 x 5?

OpenStudy (anonymous):

yep

OpenStudy (clobber02):

2/10

OpenStudy (clobber02):

or 1/5

OpenStudy (anonymous):

1/5

OpenStudy (clobber02):

yaaaasssssss

OpenStudy (anonymous):

any more?

OpenStudy (clobber02):

yes plz

OpenStudy (anonymous):

didn't we just do that?

OpenStudy (clobber02):

no not that i know of

OpenStudy (clobber02):

yup we did

OpenStudy (anonymous):

the question you asked was "2 - (-2) to the power of -2" LOL

OpenStudy (clobber02):

yup stupid me lol

OpenStudy (clobber02):

can u help me with a few more i have been sick and in the hospital and i have this worksheet/quiz i have to turn in but i am still in the hospital can u help me with 5 more plz

OpenStudy (clobber02):

?

OpenStudy (anonymous):

sure! : )

OpenStudy (clobber02):

quicker would be better because it is 10:25 here and y doc is saying i have to go to bed at 10:35

OpenStudy (anonymous):

are there any choices?

OpenStudy (anonymous):

a+c>b

OpenStudy (clobber02):

what does that mean

OpenStudy (anonymous):

side a and c has to be greater that side b

OpenStudy (clobber02):

ok

OpenStudy (clobber02):

A?

OpenStudy (anonymous):

sooo all of them are not fit for the perimeter of a triangle

OpenStudy (clobber02):

so what do i put

OpenStudy (anonymous):

all, cause they are not fit...

OpenStudy (clobber02):

it only lets me pick 1

OpenStudy (anonymous):

then 72 cause that is the highest and most likely...

OpenStudy (clobber02):

k

OpenStudy (anonymous):

next....

OpenStudy (anonymous):

maybe 11.....

OpenStudy (clobber02):

Mr. Patrick teaches math to 1515 students. He was grading tests and found that when he graded everyone's test except Payton's, the average grade for the class was 8080. After he graded Payton's test, the test average became 8181. What was Payton's score on the test?

OpenStudy (clobber02):

Mr. Patrick teaches math to 15 students. He was grading tests and found that when he graded everyone's test except Payton's, the average grade for the class was 80. After he graded Payton's test, the test average became 81. What was Payton's score on the test?

OpenStudy (anonymous):

what do you think?

OpenStudy (clobber02):

a or b

OpenStudy (clobber02):

?

OpenStudy (anonymous):

where are the choices?

OpenStudy (anonymous):

i think a.....

OpenStudy (clobber02):

thank you i have to go to bed now

OpenStudy (anonymous):

OK, cya!